# MoDA: Multi-modal Diffusion Architecture for Talking Head Generation

### v2

@@ Abstract
Talking head generation with arbitrary identities and speech audio remains a crucial problem in the realm of the virtual metaverse. Recently, diffusion models have become a popular generative technique in this field with their strong generation capabilities. However, several challenges remain for diffusion-based methods: 1) inefficient inference and visual artifacts caused by the implicit latent space of Variational Auto-Encoders (VAE), which complicates the diffusion process; 2) a lack of authentic facial expressions and head movements due to inadequate multi-modal information fusion. In this paper, MoDA handles these challenges by: 1) defining a joint parameter space that bridges motion generation and neural rendering, and leveraging flow matching to simplify diffusion learning; 2) introducing a multi-modal diffusion architecture to model the interaction among noisy motion, audio, and auxiliary conditions, enhancing overall facial expressiveness. In addition, a coarse-to-fine fusion strategy is employed to progressively integrate different modalities, ensuring effective feature fusion. Experimental results demonstrate that MoDA improves video diversity, realism, and efficiency, making it suitable for real-world applications. Project Page: https://lixinyyang.github.io/MoDA.github.io/
@@ 1 Introduction
Talking head generation creates a speaking portrait from a single image, guided by audio and other modalities. Despite advances in lip-synchronization, generating realistic and efficient videos remains a persistent challenge. The early methods [22, 25] based on Generative Adversarial Networks (GANs) [10] were the primary solution for audio-to-portrait translation, demonstrating relatively accurate lip-synchronization. However, these methods remain limited in capturing expressive emotions and natural head movements. Recently, diffusion models [7, 13, 31] have significantly advanced generative modeling, enabling the creation of highly diverse and high-fidelity images and videos. Motivated by this progress, significant research efforts [5, 3, 42, 2] have been made aimed at extending these innovations to the development of portrait image animation.
These methods are typically categorized into single-stage and two-stage frameworks. The former [5, 36, 15, 41] rely on a VAE-style latent space to directly generate video content from the audio input, integrating both motion and appearance in one pass. Despite being conceptually simple and trainable from end to end, two major limitations still persist, as shown in Fig. 2: 1) inefficient inference process. Most diffusion-based approaches rely on VAEs to model both lip motion diversity and visual appearance, but the implicit latent space and decoding complexity of VAEs significantly increase computational overhead, limiting their suitability for real-time applications; 2) limited capability for long-term video generation and visual artifacts. Methods like EMO [36] generate new video content by conditioning on previous frames. These prior frames convey motion cues but also leak appearance details, causing pixel errors to accumulate and degrade video quality.
The two-stage methods [42, 2, 17] use diffusion models to generate low-dimensional intermediate motion representations, simplifying the learning process. However, the accuracy of these intermediate representations, and therefore the final video quality, is often limited by inadequate multi-modal fusion. A key reason is that conventional cross-attention allows information to flow in just one direction, typically from audio to motion, preventing audio cues from being refined with motion context. As shown in ablation study, this limitation produces motion sequences that become inconsistent when the model is conditioned on arbitrary identities inputs.
In this paper, we propose a novel framework named MoDA, designed to generate realistic talking head videos. MoDA first establishes a joint parameter space that bridges motion generation and neural rendering, consisting of a disentangled motion-appearance space [12] along with the corresponding audio, emotion, and identity information. To address the fusion bottleneck, we draw inspiration from recent lip-to-speech tasks [38, 24], visual information can provide contextual cues complementary to audio. As shown in Fig. 3, MoDA introduces the Multi-modal Diffusion Transformer (MMDiT) [9], equipped with rectified flow [19], as the backbone diffusion framework to facilitate multi-modal fusion. In this design, MoDA can dynamically adapt audio features based on motion, identity, and emotion, thereby improving the accuracy of motion generation.
Another cause of inadequate multi-modal fusion is the frequent neglect of semantic overlaps within the joint parameter space: for example, audio inherently contains emotional and identity-related cues. Independently learned weights prevent the model from fully exploiting these intrinsic correlations, ultimately hindering its overall performance. Therefore, a coarse-to-fine fusion architecture is further proposed to progressively enhance feature integration across modalities. Initially, the model maintains modality-specific weights to capture the unique characteristics of each condition. In the intermediate stage, audio and auxiliary conditions, such as identity and emotion, begin to share attention weights. Finally, all modalities are merged into a unified representation space, allowing holistic fusion and improving the overall generative expressiveness. Moreover, to further encourage precise lip-synchronization while maintaining motion diversity, this paper provides an optional Adaptive Lip-motion Synchronization Expert (ALSE), which can be integrated during training. The contributions of MoDA can be summarized as follows:
@@ 3.1 Preliminaries
Denoising Diffusion probabilistic models (DDPMs) [32] have emerged as a powerful framework for generative modeling by formulating the data generation process as an iterative denoising procedure. In the forward diffusion process, Gaussian noise $\epsilon$ is gradually introduced into the data distribution across $T$ discrete timesteps, producing noisy latent features:$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$ where $\alpha_{t}$ represents a variance schedule that determines the noise level at each timestep, and $z_{0}$ is the raw data. The model is trained to reverse this process by taking the noisy latent representation $z_{t}$ as input and estimating the added noise $\epsilon$. The training objective is defined as:$\mathcal{L}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ where $\epsilon_{\theta}$ denotes the noise prediction generated by the model, and $c$ represents additional conditioning signals, such as audio or motion frames, which are particularly relevant in the generation of talking videos.
@@
$\mathcal{L}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||(\epsilon-z_{0})-v_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ (1)
@@ 3.2.1 Joint Parameter Space
As shown in Fig. 3, we incorporate the existing facial re-enactment framework [12] to extract disentangled facial representations. Specifically, the motion extractor $\mathcal{M}$ yields expression deformations $\delta$, head pose parameters $(R,t)$, the canonical keypoints of the source image $x_{c}$, and a scaling factor $s$. The motion representation $(R_{s},\delta_{s},t_{s},s_{s})\in\mathbb{R}^{70}$, serves as an identity-agnostic representation of the source input and is used to train MoDA to predict $(\hat{R},\hat{\delta},\hat{t},\hat{s})$ given audio input.
@@ 3.2.3 Coarse-to-Fined Feature Fusion
Although assigning modality-specific paths with separate QKV projections and FeedForward Networks (FFNs) in the attention mechanism can enhance multi-modal fusion, this design often overlook the inherent semantic commonalities shared across modalities. Independently learned weights hinder effective multi-modal feature fusion and introduce redundant parameters, potentially causing convergence issues. This issue is particularly pronounced in tasks like audio-to-motion generation, where no pixel-level inputs are involved and the semantic gap between modalities is relatively small. In such cases, maintaining entirely separate parameterizations fails to provide meaningful benefits and instead leads to duplicated representations. To enhance the model’s capacity for multi-modal understanding, we introduce a Coarse-to-Fine Feature Fusion strategy. This fusion strategy progressively integrates multi-modal features while reducing unnecessary parameters, thereby enhancing the multi-modal understanding and training stability.
As illustrated in Fig. 3, the architecture progresses through three stages: a four-stream, two-stream, and single-stream process. In the four-stream stage, each modality is processed independently to learn modality-specific representations, facilitating early-stage feature differentiation and alignment. In the two-stream stage, the audio stream is concatenated with emotion and identity features to form a merged stream that shares weights. This design encourages a balanced integration of emotional and identity cues from both the audio and the reference image. Finally, all modalities are unified into a single representation stream to allow deeper fusion and enhance the generative expressiveness.
@@ 3.2.4 Loss Function
$\mathcal{L}_{RF}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||x-v_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ (4)
@@
$\mathcal{L}_{vel}=||x^{\prime}-m^{\prime}||_{2}^{2}+||x^{\prime\prime}-m^{\prime\prime}||_{2}^{2},$ (5)
where $m$ denotes the output of $v_{\theta}(z_{t},t,c)$, and $m^{\prime\prime}$, $m^{\prime\prime}$ denote the first-order and second-order derivatives of $m$. To further enforce lip-sync accuracy, we design the ALSE pretrained in audio and motion features and compute the loss $\mathcal{L}_{ALSE}$. It is worth noting that supervision from the pretrained ALSE is optional and can be applied as needed. In summary, the final loss can be expressed as follows:
@@
$\lambda_{sync}$ =$=\begin{cases}1,&\text{if}\ \mathcal{L}_{ALSE}<\tau\\ 0,&\text{otherwise}\end{cases}$ , (7)
@@ 4.1.1 Dataset and Metrics
Our motion generation model is primarily trained on three publicly available datasets: HDTF [49], CelebV-Text [46], and MEAD [39]. For evaluation, we conducted experiments on three distinct test sets. The first two are derived from public datasets, CelebV-HQ [50] and HDTF, each consisting of 50 randomly sampled clips ranging from 3 to 10 seconds in length. The third is an in-the-wild set with 20 diverse cases, including real individuals, animated characters, dynamic scenes, and complex headwear. Each sample is paired with audio that is speech, emotional dialogue, or singing. We utilize several evaluation metrics to assess the performance of the proposed method. The Fréchet Inception Distance (FID) [28] and the Fréchet Video Distance (FVD) [37] are used to assess the quality of the generated output, while the F-SIM [36] measures facial similarity. In addition, Sync-C [4] and Sync-D [4] metrics are introduced to evaluate lip-synchronization between different methods. A temporal smoothness metric (Smo) [14] is also utilized to monitor the continuity of generated motion.

### v3

@@ Abstract
Talking head generation with arbitrary identities and speech audio remains a crucial problem in the realm of the virtual metaverse. Despite progress, current methods still struggle to synthesize diverse facial expressions and natural head movements while generating synchronized lip movements with the audio. The main challenge is stylistic discrepancies between speech audio, individual identity, and portrait dynamics. To address the challenge of inter-modal inconsistency, we introduce MoDA, a multi-modal diffusion architecture with two well-designed technologies. First, MoDA explicitly models the interaction among motion, audio, and auxiliary conditions, enhancing overall facial expressions and head dynamics. In addition, a coarse-to-fine fusion strategy is employed to progressively integrate different conditions, ensuring effective feature fusion. Experimental results demonstrate that MoDA improves video diversity, realism, and efficiency, making it suitable for real-world applications. Project Page: https://lixinyyang.github.io/MoDA.github.io/
@@ 1 Introduction
Talking head generation aims to create a photorealistic, speaking portrait from a single image, guided by audio and other modalities. Combined with the generative adversarial network (GAN) [9] and diffusion model [28], recent methods demonstrate widespread potential applications, such as immersive telepresence, virtual characters, and augmented reality.
Diffusion models have recently marked a significant advancement in generative modeling, enabling the creation of highly diverse videos. Early diffusion-based methods [5, 33, 13, 38] generate the final video directly from the audio input. Although trainable from end to end, methods like Hallo2 [5] remain two major limitations persist, as shown in Fig. 2: 1) Inefficient inference process and visual artifacts. 2) Unnatural facial expressions and head movements with precise lip-sync. Recently, two-stage methods [39, 2, 15] have simplified the diffusion process by bypassing complex variational auto-encoder (VAE) decoding. Methods like VASA-1 [39] first use the diffusion model to generate intermediate motion representations from audio, and then use a separate rendering network to synthesize the final video. However, the final video quality is heavily dependent on the accuracy of these intermediate representations. Thus, these methods still struggle to achieve natural facial dynamics with precise lip-sync due to suboptimal predictions.
Looking into the aforementioned issues, we argue that their root cause is the stylistic discrepancies between speech audio, individual identity, and portrait dynamics. These methods typically concatenate multiple conditions to form a mixed representation, which is then fed into a cross-attention mechanism where information flows only from this mixed modality to motion. This design introduces a learning bias, causing the model to focus only on the most shallow feature cues while neglecting the intricate relationships between the modalities. Consequently, it fails to handle more complex or conflicting scenarios. As in the ablation study, this limitation leads directly to inconsistent motion sequences when the model is conditioned on arbitrary identity input.
To address the challenge of inter-modal inconsistency, we propose MoDA, a novel framework designed for synergistic talking head generation. MoDA begins by operating within a joint parameter space that bridges motion generation and neural rendering, encompassing a disentangled motion-appearance space [11] along with audio, emotion, and identity. This space has a dimensionality that is an order of magnitude lower than traditional VAE spaces, which dramatically reduces the complexity of multi-modal fusion. Moreover, by incorporating optional conditions like identity and emotion, MoDA makes the generative modeling of complex distribution more tractable and increases fine-grained control over the generative process.
MoDA is guided by two core principles designed to address these inconsistencies: 1) We draw inspiration from recent lip-to-speech tasks [35, 21], where visual information can provide additional context to complement the audio. As shown in Fig. 3, MoDA introduces the Multi-modal Diffusion Transformer (MMDiT) [8], equipped with rectified flow [17], as the framework to facilitate multi-modal fusion. In this design, MoDA can dynamically adapt audio features based on motion, identity, and emotion, thereby improving the accuracy of motion generation. 2) To systematically integrate these modalities based on semantic information, MoDA implements a coarse-to-fine fusion strategy. Initially, the model uses separate weights to capture the unique characteristics of each modality. In the intermediate stage, a unified representational space is introduced for semantically linked modalities like audio, emotion, and identity to form a unified motion command. Finally, all modalities are integrated into a unified representation space, allowing holistic fusion. To further encourage precise lip-sync while maintaining motion diversity, MoDA provides an optional Adaptive Lip-motion sync Expert (ALSE), which can be integrated during training. The contributions can be summarized as follows:
@@ 3.1 Preliminaries
[…]
@@
[…]
@@ 3.2.1 Joint Parameter Space
[…]
@@ 3.2.3 Coarse-to-Fined Feature Fusion
Although assigning modality-specific paths with separate QKV projections and FeedForward Networks (FFNs) in the attention mechanism can enhance multi-modal information, this design often overlooks the inherent semantic commonalities shared across modalities. Independently learned weights hinder effective multi-modal feature fusion and introduce redundant parameters, potentially causing inconsistency issues. This issue is particularly pronounced in tasks like audio-to-motion generation, where no pixel-level inputs are involved, and the semantic gap between modalities is relatively small. In such cases, maintaining entirely separate parameterizations fails to provide meaningful benefits and instead leads to duplicated representations. To enhance the model’s capacity for multi-modal understanding, we introduce a Coarse-to-Fine Feature Fusion strategy. This fusion strategy progressively integrates multi-modal features while reducing unnecessary parameters, thereby enhancing the multi-modal understanding and training stability.
As illustrated in Fig. 3, the architecture progresses through three stages: a four-stream, two-stream, and single-stream process. In the four-stream stage, each modality is processed independently to learn modality-specific representations, facilitating early-stage feature differentiation. In the two-stream stage, the audio stream is concatenated with emotion and identity features to form a merged stream that shares weights. This design encourages a balanced integration of emotional and identity cues from both the audio and the reference image. Finally, all modalities are unified into a single representation stream to allow deeper fusion and enhance generative expressiveness.
@@ 3.2.4 Loss Function
[…]
@@
[…]
where $m$ denotes the output of $v_{\theta}(z_{t},t,c)$, and $^{\prime\prime}, m′′m^{\prime\prime}$ denote the first-order and second-order derivatives of $m$. To further enforce lip-sync accuracy, we design the ALSE pretrained in audio and motion features and compute the loss $\mathcal{L}_{ALSE}$. It is worth noting that supervision from the pretrained ALSE is optional and can be applied as needed. In summary, the final loss can be expressed as follows:
@@
$\lambda_{sync}$ $=\begin{cases}1,&\text{if}\ \mathcal{L}_{ALSE}<\tau\\ 0,&\text{otherwise}\end{cases}$ , (7)
@@ 4.1.1 Dataset and Metrics
MoDA is primarily trained on three publicly available datasets: HDTF [46], CelebV-Text [43], and MEAD [36]. For evaluation, we conducted experiments on three distinct test sets. The first two are derived from public datasets, CelebV-HQ [47] and HDTF, each consisting of 50 randomly sampled clips ranging from 3 to 10 seconds in length. The third is an in-the-wild set with 20 diverse cases, including real individuals, animated characters, dynamic scenes, and complex headwear. Each sample is paired with audio that is speech, emotional dialogue, or singing. We utilize several evaluation metrics to assess the performance of the proposed method. The Fréchet Inception Distance (FID) [25] and the Fréchet Video Distance (FVD) [34] are used to assess the quality of the generated output, while the F-SIM [33] measures facial similarity. In addition, Sync-C [4] and Sync-D [4] metrics are introduced to evaluate lip-synchronization between different methods. A temporal smoothness metric (Smo) [12] is also utilized to monitor the continuity of generated motion.